# Distributionally Robust Optimization with Adversarial Data Contamination

### v1

@@ D.3 Further Details for SVM Example
This section provides additional details for the Support Vector Machine (SVM) case discussed in Example 4.4. We consider a binary classification problem. Let the samples be $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\{\pm 1\}$, drawn from a reference distribution ${\mathbb{P}}_{0}$. The transportation cost for the Wasserstein distance is given by $c(({\bm{x}},y),({\bm{x}}^{\prime},y^{\prime}))=\chi(y-y^{\prime})+\|{\bm{x}}-{\bm{x}}^{\prime}\|_{r}$, where $\chi(0)=0$ and $\chi(z)=\infty$ for $z\neq 0$ (ensuring labels do not change during transport), and $\|\cdot\|_{r}$ is the $\ell_{r}$-norm for covariates. The loss function is the hinge loss, $\ell({\bm{w}};{\bm{x}},y)=(1-y{\bm{w}}\cdot{\bm{x}})_{+}$.
@@
As established in Section 3.3 (specifically, Equation 3), and noting that the hinge loss $(1-z)_{+}$ is $1$-Lipschitz (so $\zeta=1$), this problem is equivalent to:
@@
where $11/s+1/r=$ 1. For Example 4.4, we consider $r=2$, which implies $s=2$ (Euclidean norms).
Assumption D.2 (Distributional Assumptions on Covariates).
We assume the following for the marginal distribution of covariates ${\bm{x}}$ under ${\mathbb{P}}_{0}$, similar to [diakonikolas2019sever]:
@@
With Assumption D.2, we can show that restricting the weight vector ${\bm{w}}$ to an $\ell_{2}$-ball of radius $1/(\lambda\sigma)$ incurs at most an $O(\lambda)$ additive error in the expected hinge loss. This is formalized in the following lemma.
Lemma D.3 (Clipping of ${\bm{w}}$ for Hinge Loss).
@@
This proof adapts the argument from [diakonikolas2019sever, Lemma E.9] to the hinge loss. Let $\ell(\hat{{\bm{w}}};{\bm{x}},y)=(1-y(\hat{{\bm{w}}}\cdot{\bm{x}}))_{+}$. Fix ${\bm{w}}\in\mathbb{R}^{d}$. If $\|{\bm{w}}\|_{2}\leq 1/(\lambda\sigma)$, then ${\bm{w}}_{P}={\bm{w}}$, and the statement is trivial. Assume $\|{\bm{w}}\|_{2}>1/(\lambda\sigma)$, so $\|{\bm{w}}_{P}\|_{2}=1/(\lambda\sigma)$. Also, ${\bm{w}}_{P}=k{\bm{w}}$ for $k=1/(\lambda\sigma\|{\bm{w}}\|_{2})<1$.
@@
On the complement event $E^{c}=\{({\bm{x}},y):|{\bm{x}}\cdot{\bm{w}}_{P}|\leq 1\}$: We have $\ell({\bm{w}}_{P};{\bm{x}},y)=(1-y({\bm{w}}_{P}\cdot{\bm{x}}))_{+}$. Since $y\in\{\pm 1\}$ and $|{\bm{x}}\cdot{\bm{w}}_{P}|\leq 1$, $y({\bm{w}}_{P}\cdot{\bm{x}})$ is between $-1$ and $1$. So, $1-y({\bm{w}}_{P}\cdot{\bm{x}})$ is between 00 and $2$. Thus, $\ell({\bm{w}}_{P};{\bm{x}},y)\leq 2$. As $\ell({\bm{w}};{\bm{x}},y)\geq 0$, on $E^{c}$ we have $\ell({\bm{w}}_{P};{\bm{x}},y)\leq 2\leq 2+\ell({\bm{w}};{\bm{x}},y)$.
We now bound ${\mathbb{P}}_{0}(E^{c})$. The condition $|{\bm{x}}\cdot{\bm{w}}_{P}|\leq 1$ implies $|{\bm{x}}\cdot({\bm{w}}_{P}/\|{\bm{w}}_{P}\|_{2})|\leq 1/\|{\bm{w}}_{P}\|_{2}=\lambda\sigma$. Let ${\bm{v}}={\bm{w}}_{P}/\|{\bm{w}}_{P}\|_{2}$. Then $\|{\bm{v}}\|_{2}=1$. So, ${\mathbb{P}}_{0}(E^{c})={\mathbb{P}}_{0}(|{\bm{x}}\cdot{\bm{w}}_{P}|\leq 1)={\mathbb{P}}_{0}(|{\bm{x}}\cdot{\bm{v}}|\leq\lambda\sigma)$. By the anti-concentration assumption (Assumption D.2), ${\mathbb{P}}_{0}(E^{c})\leq B\lambda$.
@@
$=\mathbb{E}[\ell({\bm{w}};{\bm{x}},y)|E]{\mathbb{P}}_{0}(E)+2{\mathbb{P}}_{0}(E^{c})+\mathbb{E}[\ell({\bm{w}};{\bm{x}},y)|E^{c}]{\mathbb{P}}_{0}(E^{c})$
$=\mathbb{E}[\ell({\bm{w}};{\bm{x}},y)]+2{\mathbb{P}}_{0}(E^{c})$
$\leq\mathbb{E}[\ell({\bm{w}};{\bm{x}},y)]+2B\lambda.$
@@
Fact D.4 (Uniform Convergence for SVM, adapted from [diakonikolas2019sever, Lemma C.3]).
Let $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ be $N$ i.i.d. samples from a distribution ${\mathbb{P}}_{0}$ satisfying Assumption D.2. Let $S=\{({\bm{x}}_{i},y_{i}):\|{\bm{x}}_{i}\|_{2}\leq 80\sigma\sqrt{d/\epsilon}\}$. Then for a sample size $N=O(d\log(d/\lambda)/\epsilon)$, with probability at least $0.9$, it holds that $|S|\geq(1-\epsilon)N$ and for all ${\bm{w}}\in{\mathcal{B}}((\lambda\sigma)^{-1})$,
@@
To obtain the $O(\epsilon^{1/4})$ error bound mentioned in Example 4.4, we set the parameter $\lambda$ from the anti-concentration assumption appropriately. The error from clipping ${\bm{w}}$ (Lemma D.3) is $O(B\lambda)$. The uniform convergence error (Fact D.4), if $\lambda\ll 1$, is $O(\lambda^{-1}\sqrt{\epsilon})$. Our main algorithm is applied to the set $S$, which can be seen as being $2\epsilon$-corrupted if the original data had $\epsilon$-corruptions and we discard an additional $\epsilon$ fraction based on norm. The error from our main algorithm (Proposition 4.1) is $O(\sigma\zeta\sqrt{2\epsilon}W)=O(\sigma\sqrt{\epsilon}(\lambda\sigma)^{-1})=O(\lambda^{-1}\sqrt{\epsilon})$ since $\zeta=1$ for hinge loss. Total error from these sources is $O(B\lambda+\lambda^{-1}\sqrt{\epsilon})$.
We want to balance these errors. To minimize this, we can set $\lambda=\lambda^{-1}\sqrt{\epsilon}$, which implies $\lambda^{2}=\sqrt{\epsilon}$, so $\lambda=\epsilon^{1/4}$. Plugging this $\lambda$ back, the error becomes $O(B\epsilon^{1/4})$. This shows that an overall error of $O(\epsilon^{1/4})$ in the objective function value is achievable viewing $B$ as an absolute constant. The radius of the weight ball is $W=1/(\lambda\sigma)=1/(\sigma\epsilon^{1/4})$.
Appendix E Tuning Step Sizes for Algorithm 1
The primary algorithm, Algorithm 1, and its theoretical guarantee in Proposition 4.1 depend on the parameter $\gamma=\frac{\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}}{\zeta\sqrt{N}}$. Specifically, $\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}$ (denoted by $D_{0}$ hereafter) is the distance from the initial iterate to the optimal solution ${\bm{w}}^{*}$ of the empirical stable objective, which is unknown in practice. This section details a practical and theoretically sound procedure for tuning $\gamma$ by performing a geometric search for $D_{0}$. While a simple search across a predefined range is possible, we propose a more efficient method that allows for early stopping, avoiding unnecessary computation if the optimal parameter setting is found long before the search range is exhausted.
A Bounded Geometric Search with Early Stopping
@@
    Run Algorithm 1 with $\gamma_{j}=D_{j}/(\zeta\sqrt{N})$ to obtain a solution $\hat{{\bm{w}}}_{T}^{(j)}$.
@@
Robust Estimation of Objective Function Value
For each candidate solution $\hat{{\bm{w}}}_{T}^{(j)}$, we need to estimate the objective function value $F_{{\mathbb{P}}_{0}}(\hat{{\bm{w}}}_{T}^{(j)})=\mathbb{E}_{({\bm{x}},y)\sim{\mathbb{P}}_{0}}[\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{x}},y)]+\psi(\hat{{\bm{w}}}_{T}^{(j)})$, where $\psi({\bm{w}})=\rho\zeta\|{\bm{w}}\|_{s}$. The regularization term $\psi(\hat{{\bm{w}}}_{T}^{(j)})$ can be computed directly. The challenge is to estimate the expected loss $\mathbb{E}_{({\bm{x}},y)\sim{\mathbb{P}}_{0}}[\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{x}},y)]$ using the $\epsilon$-corrupted training samples $\{({\bm{x}}_{i}^{c},y_{i}^{c})\}_{i=1}^{N}$.
We compute the loss values $L_{i}^{(j)}=\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{x}}_{i}^{c},y_{i}^{c})$ for $i=1,\dots,N$. This set $\{L_{i}^{(j)}\}_{i=1}^{N}$ constitutes an $\epsilon$-corrupted sample of the random variable $L^{(j)}({\bm{\xi}})=\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{\xi}})$, where ${\bm{\xi}}\sim{\mathbb{P}}_{0}$. We can robustly estimate the mean of $L^{(j)}({\bm{\xi}})$ from these corrupted samples. Standard robust mean estimation techniques for 1-dimensional data (such as those based on truncated mean [diakonikolas2022algorithmic, Section 1.4.2] or variants of Algorithm 3 adapted to 1-dimensional settings) achieve an estimation error of $O(\sigma_{L^{(j)}}\sqrt{\epsilon})$; see Fact 2.7, where $\sigma_{L^{(j)}}$ is the standard deviation of the clean loss values $L^{(j)}({\bm{\xi}})$.
To bound $\sigma_{L^{(j)}}$: The loss function $\ell({\bm{w}};{\bm{x}},y)=\phi({\bm{w}}\cdot{\bm{x}},y)$ has $\phi(\cdot,y)$ as $\zeta$-Lipschitz. The iterates $\hat{{\bm{w}}}_{T}^{(j)}$ are bounded. By Corollary D.1, $\|\hat{{\bm{w}}}_{T}^{(j)}\|_{2}=O(\|{\bm{w}}_{0}\|_{2}+\|{\bm{w}}^{*}\|_{2})$. By Assumption 3.4, the arguments ${\bm{w}}\cdot{\bm{x}}$ will have variance $\|{\bm{w}}\|_{2}^{2}\sigma^{2}$. The Lipschitz property of the loss $\ell_{y}$ implies that the standard deviation $\sigma_{L^{(j)}}$ of the loss values is $O(\zeta\|\hat{{\bm{w}}}_{T}^{(j)}\|_{2}\sigma)$. Consequently, the error in estimating $\mathbb{E}_{({\bm{x}},y)\sim{\mathbb{P}}_{0}}[\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{x}},y)]$ is $E_{\text{eval}}=O(\zeta(\|{\bm{w}}_{0}\|_{2}+\|{\bm{w}}^{*}\|_{2})\sigma\sqrt{\epsilon})$.
Early Stopping Condition
@@
Stop at step $\text{Stop at step }k\text{ if }\hat{F}^{(k)}>\hat{F}_{\text{best}}+3E_{\text{ub}}\;,$ (20)
@@
Overall Error Guarantee and Complexity

### v2

@@ 4.1 Application to SVMs: Removing Dependence on the Optimal Solution Norm
We consider a binary classification problem. Let the samples be $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\{\pm 1\}$, drawn from a reference distribution ${\mathbb{P}}_{0}$. The transportation cost for the Wasserstein distance is given by $c(({\bm{x}},y),({\bm{x}}^{\prime},y^{\prime}))=\chi(y-y^{\prime})+\|{\bm{x}}-{\bm{x}}^{\prime}\|_{r}$, where $\chi(0)=0$ and $\chi(z)=\infty$ for $z\neq 0$ (ensuring labels do not change during transport), and $\|\cdot\|_{r}$ is the $\ell_{r}$-norm for covariates. The loss function is the hinge loss, $\ell({\bm{w}};{\bm{x}},y)=(1-y{\bm{w}}\cdot{\bm{x}})_{+}$.
@@
As established in Section˜3.3 (specifically, Equation˜6), and noting that the hinge loss $(1-z)_{+}$ is $1$-Lipschitz (so $\zeta=1$), this problem is equivalent to:
@@
where $1/s+1/r=1$. In the remaining subsection, we consider $r=2$, which implies $s=2$ (Euclidean norms).
Assumption 4.9 (Distributional Assumptions on Covariates).
We assume the following for the marginal distribution of covariates $\bm{x}$} under ${\mathbb{P}}_{0}$, similar to [Dia+19a]:
@@
With Assumption˜4.9, we can show that restricting the weight vector $\bm{w}$} to an $\ell_{2}$-ball of radius $1/(\lambda\sigma)$ incurs at most an $O(\lambda)$ additive error in the expected hinge loss. This is formalized in the following lemma.
Lemma 4.10 (Clipping of $\bm{w}$} for Hinge Loss).
@@
This proof adapts the argument from [Dia+19a, Lemma E.9 ] to the hinge loss. Let $\ell(\hat{{\bm{w}}};{\bm{x}},y)=(1-y(\hat{{\bm{w}}}\cdot{\bm{x}}))_{+}$. Fix ${\bm{w}}\in\mathbb{R}^{d}$. If $\|{\bm{w}}\|_{2}\leq 1/(\lambda\sigma)$, then ${\bm{w}}_{P}={\bm{w}}$, and the statement is trivial. Assume $\|{\bm{w}}\|_{2}>1/(\lambda\sigma)$, so $\|{\bm{w}}_{P}\|_{2}=1/(\lambda\sigma)$. Also, ${\bm{w}}_{P}=k{\bm{w}}$ for $k=1/(\lambda\sigma\|{\bm{w}}\|_{2})<1$.
@@
On the complement event $E^{c}=\{({\bm{x}},y):|{\bm{x}}\cdot{\bm{w}}_{P}|\leq 1\}$: We have $\ell({\bm{w}}_{P};{\bm{x}},y)=(1-y({\bm{w}}_{P}\cdot{\bm{x}}))_{+}$. Since $y\in\{\pm 1\}$ and $|{\bm{x}}\cdot{\bm{w}}_{P}|\leq 1$, $y({\bm{w}}_{P}\cdot{\bm{x}})$ is between 1-1 and $1$. So, $1-y({\bm{w}}_{P}\cdot{\bm{x}})$ is between 0 and $2$. Thus, $\ell({\bm{w}}_{P};{\bm{x}},y)\leq 2$. As $\ell({\bm{w}};{\bm{x}},y)\geq 0$, on $E^{c}$ we have $\ell({\bm{w}}_{P};{\bm{x}},y)\leq 2\leq 2+\ell({\bm{w}};{\bm{x}},y)$.
We now bound ${\mathbb{P}}_{0}(E^{c})$. The condition $|{\bm{x}}\cdot{\bm{w}}_{P}|\leq 1$ implies $|{\bm{x}}\cdot({\bm{w}}_{P}/\|{\bm{w}}_{P}\|_{2})|\leq 1/\|{\bm{w}}_{P}\|_{2}=\lambda\sigma$. Let ${\bm{v}}={\bm{w}}_{P}/\|{\bm{w}}_{P}\|_{2}$. Then $\|{\bm{v}}\|_{2}=1$. So, ${\mathbb{P}}_{0}(E^{c})={\mathbb{P}}_{0}(|{\bm{x}}\cdot{\bm{w}}_{P}|\leq 1)={\mathbb{P}}_{0}(|{\bm{x}}\cdot{\bm{v}}|\leq\lambda\sigma)$. By the anti-concentration assumption (Assumption˜4.9), ${\mathbb{P}}_{0}(E^{c})\leq B\lambda$.
@@
$=\mathbb{E}[\ell({\bm{w}};{\bm{x}},y)]+2{\mathbb{P}}_{0}(E^{c})\leq\mathbb{E}[\ell({\bm{w}};{\bm{x}},y)]+2B\lambda.$
@@
Fact 4.11 (Uniform Convergence for SVM, adapted from [Dia+19a, Lemma C.3 ]).
Let the set $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ be $N$ i.i.d. samples from a distribution ${\mathbb{P}}_{0}$ satisfying Assumption˜4.9. Let $S=\{({\bm{x}}_{i},y_{i}):\|{\bm{x}}_{i}\|_{2}\leq 80\sigma\sqrt{d/\epsilon}\}$. Then for a sample size $N=O(d\log(d/\lambda)/\epsilon)$, with probability at least $0.9$, it holds that $|S|\geq(1-\epsilon)N$ and for all ${\bm{w}}\in{\mathcal{B}}((\lambda\sigma)^{-1})$,
@@
To obtain the $O(\epsilon^{1/4})$ error bound, we set the parameter $\lambda$ from the anti-concentration assumption appropriately. The error from clipping $\bm{w}$} (Lemma 4.10) is $O(B\lambda)$. The uniform convergence error (Fact 4.11), if $\lambda\ll 1$, is $O(\lambda^{-1}\sqrt{\epsilon})$. Our main algorithm is applied to the set $S$, which can be seen as being $2\epsilon$-corrupted if the original data had $\epsilon$-corruptions and we discard an additional $\epsilon$ fraction based on norm. The error from our main algorithm (Proposition˜4.1) is $O(\sigma\zeta\sqrt{2\epsilon}W)=O(\sigma\sqrt{\epsilon}(\lambda\sigma)^{-1})=O(\lambda^{-1}\sqrt{\epsilon})$ since $\zeta=1$ for hinge loss. Total error from these sources is $O(B\lambda+\lambda^{-1}\sqrt{\epsilon})$.
We want to balance these errors. To minimize this quantity, we can set $\lambda=\lambda^{-1}\sqrt{\epsilon}$, which implies $\lambda^{2}=\sqrt{\epsilon}$, so $\lambda=\epsilon^{1/4}$. Plugging this $\lambda$ back, the error becomes $O(B\epsilon^{1/4})$. This shows that an overall error of $O(\epsilon^{1/4})$ in the objective function value is achievable viewing $B$ as an absolute constant. The radius of the weight ball is $W=1/(\lambda\sigma)=1/(\sigma\epsilon^{1/4})$.
4.2 Tuning the Step Size Parameter
A practical issue with Algorithm˜1 is that its step size parameter $\gamma$ depends on the unknown quantity $\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}$ (denoted by $D_{0}$ hereafter). Fortunately, this is not a significant barrier. We can tune $\gamma$ using a standard geometric search for this distance, which finds a sufficiently good estimate by running the algorithm a logarithmic number of times. This section details this procedure, including a robust method for evaluating candidate solutions and an optional early stopping criterion. The overhead for this tuning is a logarithmic factor in the runtime and a small additive term in the final error bound.
4.2.1 A Bounded Geometric Search with Early Stopping
@@
    Run Algorithm˜1 with $\gamma_{j}=D_{j}/(\zeta\sqrt{N})$ to obtain a solution $\hat{{\bm{w}}}_{T}^{(j)}$.
@@
4.2.2 Robust Estimation of Objective Function Value
For each candidate solution $\hat{{\bm{w}}}_{T}^{(j)}$, we need to estimate the objective function value
$F_{{\mathbb{P}}_{0}}(\hat{{\bm{w}}}_{T}^{(j)})=\mathbb{E}_{({\bm{x}},y)\sim{\mathbb{P}}_{0}}[\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{x}},y)]+\psi(\hat{{\bm{w}}}_{T}^{(j)}),$
where $\psi({\bm{w}})=\rho\zeta\|{\bm{w}}\|_{s}$. The regularization term $\psi(\hat{{\bm{w}}}_{T}^{(j)})$ can be computed directly. The challenge is to estimate the expected loss $\mathbb{E}_{({\bm{x}},y)\sim{\mathbb{P}}_{0}}[\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{x}},y)]$ using the $\epsilon$-corrupted training samples $\{({\bm{x}}_{i}^{c},y_{i}^{c})\}_{i=1}^{N}$.
We compute the loss values $L_{i}^{(j)}=\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{x}}_{i}^{c},y_{i}^{c})$ for $i=1,\dots,N$. This set $\{L_{i}^{(j)}\}_{i=1}^{N}$ constitutes an $\epsilon$-corrupted sample of the random variable $L^{(j)}({\bm{\xi}})=\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{\xi}})$, where ${\bm{\xi}}\sim{\mathbb{P}}_{0}$. We can robustly estimate the mean of $L^{(j)}({\bm{\xi}})$ from these corrupted samples. Standard robust mean estimation techniques for 1-dimensional data (such as those based on truncated mean [DK23a, Section 1.4.2] or variants of Algorithm˜3 adapted to 1-dimensional settings) achieve an estimation error of $O(\sigma_{L^{(j)}}\sqrt{\epsilon})$; see Fact˜2.7, where $\sigma_{L^{(j)}}$ is the standard deviation of the clean loss values $L^{(j)}({\bm{\xi}})$.
To bound $\sigma_{L^{(j)}}$, observe that the loss function $\ell({\bm{w}};{\bm{x}},y)=\phi({\bm{w}}\cdot{\bm{x}},y)$ has $\phi(\cdot,y)$ as $\zeta$-Lipschitz. The iterates $\hat{{\bm{w}}}_{T}^{(j)}$ are bounded. By Corollary˜4.3, $\|\hat{{\bm{w}}}_{T}^{(j)}\|_{2}=O(\|{\bm{w}}_{0}\|_{2}+\|{\bm{w}}^{*}\|_{2})$. By Assumption˜3.8, the arguments ${\bm{w}}\cdot{\bm{x}}$ will have variance $\|{\bm{w}}\|_{2}^{2}\sigma^{2}$. The Lipschitz property of the loss $\ell_{y}$ implies that the standard deviation $\sigma_{L^{(j)}}$ of the loss values is $O(\zeta\|\hat{{\bm{w}}}_{T}^{(j)}\|_{2}\sigma)$. Consequently, the error in estimating $\mathbb{E}_{({\bm{x}},y)\sim{\mathbb{P}}_{0}}[\ell(\hat{{\bm{w}}}_{T}^{(j)};{\bm{x}},y)]$ is $E_{\text{eval}}=O(\zeta(\|{\bm{w}}_{0}\|_{2}+\|{\bm{w}}^{*}\|_{2})\sigma\sqrt{\epsilon})$.
4.2.3 Early Stopping Condition
@@
$\text{Stop at step }k\text{ if }\hat{F}^{(k)}>\hat{F}_{\text{best}}+3E_{\text{ub}}\;,$ (15)
@@
4.2.4 Overall Error Guarantee and Complexity